OpenStudy (anonymous):

Find the range of the function. f(x) = (x + 5)2 + 8 y > -8 y ≥ 8 y > 8 All real numbers

OpenStudy (anonymous):

What do you think it is?

OpenStudy (luigi0210):

Welcome to openstudy :D

OpenStudy (jdoe0001):

what's the range of the function? the value of "y"? the value of "x"?

OpenStudy (anonymous):

I know it's not C I just need help ;-;

OpenStudy (anonymous):

Thanks @Luigi0210

OpenStudy (anonymous):

:3

OpenStudy (luigi0210):

By range it means what numbers can you plug into the equation to get an output.

OpenStudy (anonymous):

I've never understood how to work the equation ;-; @Luigi0210

OpenStudy (anonymous):

I doubt it's the first one then.. Is it B maybe?

OpenStudy (anonymous):

Everyone left me didn't they? .-.

OpenStudy (luigi0210):

Yup

OpenStudy (luigi0210):

I was thinking B too

OpenStudy (anonymous):

Thanks :3

OpenStudy (luigi0210):

Mhm :)

OpenStudy (anonymous):

minimum value of (x+5)^2 is zero, so minimum value of f(x)=0+8=8 maximum value of \[f \left( x \right) is \infty hence range is \ge 8\]